If an x-ray beam is incident on a 3mm thick sample and 50% of the beam is transmitted what is the average absorption coefficient of the material ?
1
Expert's answer
2019-06-19T10:23:29-0400
"e^{-\mu x}=0.5"

"\mu=\frac{\ln{2}}{x}"

"\mu=\frac{\ln{2}}{0.003}=231\ m^{-1}"